OpenStudy (anonymous):

[9.06] Tabitha wants to hang a painting in a gallery but the painting and frame must have an area of 58 square feet. The painting is 7 feet wide by 8 feet long. The frame must have an equal width of x on each side. Which quadratic equation represents the area of the painting and frame combined? x2 + 15x − 2 = 0 x2 + 15x + 58 = 0 4x2 + 30x − 2 = 0 4x2 + 30x + 58 = 0

OpenStudy (anonymous):

option a

OpenStudy (anonymous):

how??? @ash2007ray

OpenStudy (anonymous):

frame width=x NOW NEW LENGHT=7+x new width=8+x area=length*width=(7+x)*(8+x) now this should be = 58 (7+x)*(8+x)=58 now expand it get the answer!!!